# Learning to Control Unknown Strongly Monotone Games

### v2

@@ 4.1 Non-expansive Mapping
The above three lemmas are sufficient for proving Theorem 1 (see Appendix). We use the properties of two time-scale stochastic approximation to show that $(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\to(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})$. This allows us to analyze the slower iteration, i.e., the control updates (8) in the limit $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$. We rewrite the iteration as a fixed point iteration with the mapping $g(\cdot)$. Since $g(\cdot)$ is non-expansive, this iteration is a version of the inexact KM algorithm [61]. We study its ODE limit and show convergence to the set of fixed points.
@@ 4.2 Finite Time Convergence Guarantees
In this subsection we detail how to obtain our finite time convergence guarantees. This result also uses the non-expansiveness on $g$, but requires a more nuanced argument that bounds the error as follows
@@
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that for all $t>T_{0}$:
@@
For $\eta_{t}=(t+T_{1})^{-\eta}$ with $10<\eta<$ 1, this recursion yields [12]
@@
The next lemma bounds the mean squared distance between $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and $\boldsymbol{\ell^{*}}$. Note that we need to bound $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ as $g(\boldsymbol{\alpha})-\boldsymbol{\alpha}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$. We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) KM algorithm. If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [63]. Since this assumption is not true in our case, we use the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
@@
Using our assumption that $\epsilon_{t}$ goes to zero, we have that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$. For the rest of the proof, we pick $T_{0}$ to be large enough such that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$, $\epsilon_{t}<\gamma/2$ and $\eta_{t}<C_{2}$ for all $t>T_{0}$, for appropriate $C_{1}$ and $C_{2}$. Returning to (4.2), for all $t>T_{0}$, we have
@@
$\leq-\|p_{t+1}-p_{t}\|^{2}.$
Here the final inequality follows from the non-expansive nature of $g(\cdot)$. This gives us
@@
Using the assumption that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$, this gives us $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}+2\epsilon_{t}^{\prime}\|q_{t}\|^{2}.$ Taking expectation and using Lemma 7, this results in the bound
@@
Here $c_{9}$ and $c_{10}$ are constants and $T_{3}=\max\{T_{1},T_{2}\}$. The final inequality follows from the fact that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ and hence $2\epsilon\geq 3\eta$. Since $\eta_{t}$ is square summable, $\eta>0.5$ and hence $14\eta-1>$ 1, so the summation is bounded by a constant. Hence all terms in the right-hand side of (4.2) are bounded by a constant. Then for some $c_{11}>0$, we have
@@ 5 Simulation Results
Figure 2: Weighted Load Balancing in Resource Allocation Games
@@
In this section, we simulate the two applications discussed in Section 2.1. In all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$. We add Gaussian noise with mean 00 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player. We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small. In each realization, $\alpha_{0}^{k}$ and $x_{n,0}^{k}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
We pick a resource allocation game (Figure 2) inspired by DSM in electricity grid [64]. We keep $N=1000$ and $d=24$. We choose the action set for each player $n$ as $\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$ Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$, and $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$. We use the reward function $r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$ Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$. We also defined $s_{i}=\sum_{n=1}^{N}a_{n,i}^{(i)}x_{n}^{i}$, i.e., the weighted load at resource $i$. The coefficients $a_{n,i}^{(i)}$ are i.i.d. samples from a Gaussian distribution with mean $1$ and variance $0.2$. The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are i.i.d. samples from a Gaussian distribution with mean $35$ and variance $5$.
@@
In Fig. 2(a), we plot the weighted load for each resource along with their target loads. We can observe that the system converges to a NE where the loads are balanced. In Fig. 2(b), we plot the weighted load for the first 3 resources. To study the effectiveness of our algorithm, we also plot the loads in the case of a uncontrolled system., i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$. It can be observed that if the system is uncontrolled, some resources are under-utilized while others are over-utilized.
@@
For the quadratic global objective (Figure 3), we consider a quadratic game. In a quadratic game, the reward of each player is of the form $r_{n}\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{Q_{n}}\boldsymbol{x}+\boldsymbol{c_{n}}\boldsymbol{x}_{n}$ for some negative definite matrix $\boldsymbol{Q_{n}}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{c_{n}}\in\mathbb{R}^{d}$.
@@
We work with $N=100$ players and $d=5$, and action sets $\{\boldsymbol{x}_{n}\mid\sum_{i=1}^{d}x_{n}^{i}\leq 10\}$ for each player $n$. Recall that the global cost is given by $\boldsymbol{x}^{\top}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. To generate $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a standard Gaussian distribution. We use $\boldsymbol{G}=\boldsymbol{H_{0}}\boldsymbol{H_{0}}^{T}$. A random $\boldsymbol{\rho}$ is obtained with each element sampled from a Gaussian distribution with mean 0 and variance $0.5$. The above matrices $\boldsymbol{Q_{n}}$ are obtained by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$. The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way as $\boldsymbol{\rho}$.
In Fig. 3(a), we observe that our algorithm is able to converge to a lower quadratic cost as compared to a system which is left uncontrolled. In Fig. 3(b), we plot the sum of rewards for our algorithm. This is compared with the sum of rewards if the global objective was directly optimized for, i.e., each player updates their action based on $\nabla\Phi(\boldsymbol{x})$ where $\Phi(\boldsymbol{x})$ is the quadratic cost. Our algorithm provides a much higher sum of rewards after convergence. The two plots together show that our algorithm can be used as a method to optimize for a global cost while still trying to maximize the rewards for players.
@@ 6 Conclusion
We studied how to control strongly monotone games where the game structure is unknown to the manager. The goal of the controller is to ensure the game satisfies target linear constraints at NE. Our algorithm only requires the manager to observe the instantaneous constraint violation. We prove that our simple online algorithm guarantees convergence to a point $\boldsymbol{x^{*}}$ which satisfies the linear constraints $\boldsymbol{A}\boldsymbol{x^{*}}=\boldsymbol{l}^{*}$ by adjusting control inputs $\boldsymbol{\alpha}$ in real-time. Further, we show that the convergence rate for our algorithm is $O(t^{-1/4})$. Our algorithm is a two time-scale algorithm, where the algorithm does not wait for the players to converge to a NE but uses the feedback from every turn. This is an encouraging result since our algorithm is easy to implement and maintains the users’ privacy. The manager does not observe individual actions, and players do not need to reveal their reward functions, which might even be unknown to them.
@@
Controlling unknown games with bandit feedback is a new online learning paradigm. The game to be controlled is a non-stationary but highly structured bandit. Viewing game control as a bandit problem can result in new exciting research questions, and new powerful game control techniques. Generalizing our game control setting beyond linear constraints can help tackle a wider class of problems.

### v3

@@ 4.1 Non-expansive Mapping
Lemmas 4-6 are sufficient to prove Theorem 1 (see the Appendix). We use two time-scale stochastic approximation to show that $(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\to(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})$. Then we can analyze the slower iteration of the control updates (8) in the limit $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$. We rewrite (8) as a fixed point iteration with the mapping $g(\cdot)$. Since $g(\cdot)$ is non-expansive, this iteration is a version of the inexact KM algorithm [56]. We study its ODE limit and show convergence to the set of fixed points.
@@ 4.2 Finite Time Convergence Guarantees
Our finite time convergence guarantees also use the non-expansiveness on $g$, but require a more nuanced argument that bounds the error as follows
@@
Under Assumptions 1-5, there exist $C^{\prime}>0$ and $T_{0}$ such that for all $t>T_{0}$:
@@
For $\eta_{t}=(t+T_{1})^{-\eta}$ with $0<\eta<1$, this yields [8]
@@
[…]
@@
Using our assumption that $\epsilon_{t}$ goes to zero, we have that $0<\epsilon_{t}^{\prime}<1/2$ for all $t>T_{0}$. For the rest of the proof, we pick $T_{0}$ to be large enough such that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$, $\epsilon_{t}<\gamma/2$ and $\eta_{t}<C_{2}$ for all $t>T_{0}$, for appropriate $C_{1}$ and $C_{2}$. Returning to (4.2), for all $t>T_{0}$, we have
@@
$\leq-\|p_{t+1}-p_{t}\|^{2},$
where the last inequality uses that $g(\cdot)$ is non-expansive. Then,
@@
Using the assumption that $0<\epsilon_{t}^{\prime}<1/2$ for all $t>T_{0}$, this gives us $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}+2\epsilon_{t}^{\prime}\|q_{t}\|^{2}.$ Taking expectation and using Lemma 7, this results in the bound
@@
Here $c_{9}$ and $c_{10}$ are constants and $T_{3}=\max\{T_{1},T_{2}\}$. The final inequality follows from the fact that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ and hence $2\epsilon\geq 3\eta$. Since $\eta_{t}$ is square summable, $\eta>0.5$ and hence $4\eta-1>1$, so the summation is bounded by a constant. Hence all terms in the right-hand side of (4.2) are bounded by a constant. Then for some $c_{11}>0$, we have
@@ 5 Simulation Results
(c) Load balancing for time-varying loads
Figure 2: Load Balancing in Resource Allocation Games
@@
(c) Cost when no $\boldsymbol{x}\in\mathcal{X}$ satisfies $2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$
@@
In this section, we simulate the two applications discussed in Section 2.1. In all experiments, we use the step-size sequences $\eta_{t}=\frac{1}{\left(t+1\right)^{0.501}}$ and $\epsilon_{t}=\frac{1}{\left(t+1\right)^{0.753}}$ since they approximately give the best rate (see Theorem 2). To model imperfect gradient information (see Assumption 3), we add Gaussian noise with mean 0 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player. We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small. In each realization, $\alpha_{0}^{k}$ and $x_{n,0}^{k}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
We simulated a demand-side management game (Figure 2) [60] with $N=1000$ and $d=24$ hours. The action sets were $\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}$, and $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $\hat{\mu}^{i}$ and variance $0.5$. Here $\hat{\mu}^{i}$ denote the typical hourly energy consumption limit of a residential customer as given in [60, Fig. 1]. Similarly, $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $\hat{\mu}$ and variance $1$, where $\hat{\mu}$ is the typical daily consumption limit. Let $s_{i}=\sum_{n=1}^{N}x_{n}^{i}$ be the energy consumption at the $i$-th hour. We use the reward function $r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$ Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $\hat{\omega}^{i}$ and variance $0.5$. The value of energy at hour $i$, $\hat{\omega}^{i}$, was adapted from the demand characteristics of residential consumers [61]. The objectives $s_{i}=\ell_{i}^{*}$ were chosen to shave the peak of the uncontrolled system [60], with the following target loads:
$\boldsymbol{\ell}^{*}=[$ $53.58,45.09,37.63,34.32,41.16,28.95,37.55,34.18,$
$27.99,28.15,28.71,28.37,19.96,27.80,23.83,27.93,$
$20.45,21.44,20.86,28.67,20.89,40.12,37.64,34.32].$
@@
In Fig. 2(a), we plot the weighted load for each resource along with their target loads. We can observe that the system converges to a NE where the loads are balanced. In Fig. 2(b), we plot the weighted load for the first 3 resources. To study the effectiveness of our algorithm, we also plot the loads in the case of an uncontrolled system, i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$. It can be observed that if the system is uncontrolled, some resources are under-utilized while others are over-utilized.
@@
For the quadratic global objective (Figure 3), we consider a quadratic game. In a quadratic game, the reward of each player is of the form $r_{n}\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{Q_{n}}\boldsymbol{x}+\boldsymbol{c_{n}}^{T}\boldsymbol{x}_{n}$ for some negative definite matrix $\boldsymbol{Q_{n}}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{c_{n}}\in\mathbb{R}^{d}$.
@@
We work with $N=100$ players and $d=5$, and action sets $\{\boldsymbol{x}_{n}\mid 0\leq\sum_{i=1}^{d}x_{n}^{i}\leq 10\}$ for each player $n$. Recall that the global cost is given by $\boldsymbol{x}^{\top}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. To generate $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a standard Gaussian distribution. We use $\boldsymbol{G}=\boldsymbol{H_{0}}\boldsymbol{H_{0}}^{T}$. A random $\boldsymbol{\rho}$ is obtained with each element sampled from a Gaussian distribution with mean 0 and variance $0.5$. The above matrices $\boldsymbol{Q_{n}}$ are obtained by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$. The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way to $\boldsymbol{\rho}$.
Fig. 3(a) shows that our algorithm converges to the optimal global cost, which is significantly better than the global cost of the uncontrolled system. Even more, Fig. 3(b) shows an advantage of our scheme over a centralized scheme that optimizes the global cost $\Phi(\boldsymbol{x})$ directly, i.e., where each player updates their action using $\nabla\Phi(\boldsymbol{x})$. While both approaches minimize the global cost, our algorithm results in a significantly better sum of rewards $\sum_{n=1}^{N}r_{n}(\boldsymbol{x})$. The two plots together show that our algorithm can optimize the global cost under the soft constraint of maintaining a high sum of local player rewards.
In Figures 2(c) and 3(c), we study the performance of our algorithm beyond our theoretical model and guarantees. In Fig. 2(c), the target loads $\boldsymbol{\ell^{*}}_{t}$ vary over time. The objective now is $\sum_{n=1}^{N}x_{n,t}^{i}=\ell_{i,t}^{*}$. Our algorithm easily tracks the time-varying loads under both smoothly varying loads and abruptly changing load profiles. In Fig. 3(c), we study the case where the linear constraints are infeasible, i.e., $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ is not satisfied by any $x\in\mathcal{X}$. In this case, we make a slight modification to our algorithm by projecting the control input $\boldsymbol{\alpha}$ onto a sphere after each iteration. This is required to ensure the stability of the algorithm. Although it is no longer possible to achieve the optimal cost, our modified algorithm still demonstrates stable behavior and convergence to an action profile with a lower quadratic cost as compared to the uncontrolled system.
@@ 6 Conclusion
We studied how to control a strongly monotone game that is unknown to the manager. The goal of the manager is to ensure the game satisfies target linear constraints at NE. Our simple algorithm only requires the manager to observe the instantaneous constraint violation. We prove that our algorithm guarantees convergence to an $\boldsymbol{x^{*}}$ which satisfies $\boldsymbol{A}\boldsymbol{x^{*}}=\boldsymbol{\ell^{*}}$ by adjusting control inputs $\boldsymbol{\alpha}$ in real-time. Further, we show that the convergence rate for our algorithm is $O(t^{-1/4})$. Our two time-scale algorithm does not wait for the players to converge to a NE but uses the feedback from every turn. This is an encouraging result since our algorithm is easy to implement and maintains the users’ privacy. The manager does not observe individual actions, and players do not reveal their reward functions, which might even be unknown to them.
@@
Controlling unknown games with bandit feedback is a new online learning paradigm. The game to be controlled is a non-stationary but highly structured bandit. Viewing game control as a bandit problem can result in new, powerful game control techniques. Generalizing our game control setting beyond linear constraints can help tackle a wider class of problems.
[…]
7 Proofs from Section 3
7.1 Proof for Theorem 1
For a fixed $\boldsymbol{\alpha}$, the ODE limit of iteration (7) is $\dot{\boldsymbol{x}}(t)=F(\boldsymbol{x}(t),\boldsymbol{\alpha})+\mathfrak{b}_{\mathcal{X}}(\boldsymbol{x}(t)).$ Here $\mathfrak{b}_{\mathcal{X}}(\boldsymbol{x}(t))$ is the boundary projection term required to keep $\boldsymbol{x}(t)$ inside $\mathcal{X}$ at all terms. Since the game is strongly monotone for any $\boldsymbol{\alpha}$, [4, Theorem 9] shows that for a fixed $\boldsymbol{\alpha}$, there exists a unique globally stable equilibrium of the above ODE, which is the NE $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$. Additionally, we have shown that $\|\boldsymbol{\alpha}_{t}\|$ has bounded expectation in Lemma 6 for all $t$, which implies that with probability 1, $\|\boldsymbol{\alpha}_{t}\|<\infty$ for all $t$. Then Part (a) of Lemma 4 shows that $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is Lipschitz continuous in $\boldsymbol{\alpha}$. Then [58, Lemma 8.1] shows that $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$. Now, we can write iteration (8) as
$\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t})+\epsilon_{t}^{\prime}\gamma\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})),$
where $\epsilon_{t}^{\prime}=\epsilon_{t}/\gamma$. The ODE limit for this iteration is
$\dot{\boldsymbol{\alpha}}(t)=g(\boldsymbol{\alpha}(t))-\boldsymbol{\alpha}(t).$ (22)
The set of fixed points of $g(\cdot)$ is $\mathcal{N}_{opt}$, which is non-empty as shown in Lemma 3. Let $\boldsymbol{\alpha^{*}}$ be any such point. [62, Theorem 3.1] shows that the solutions of $\dot{\boldsymbol{y}}(t)=f(\boldsymbol{y}(t))-\boldsymbol{y}(t)$ converge to the set of fixed points of $f(\cdot)$, when $f$ is non-expansive under $\ell_{2}$ norm and the set of fixed points of $f(\cdot)$ is non-empty. Since $\boldsymbol{\alpha}(t)$ converges to a point in $\mathcal{N}_{opt}$, this completes our proof that the iterates $\boldsymbol{\alpha}_{t}$ converge to $\mathcal{N}_{opt}$.
7.2 Proof for Theorem 2
Let $T_{0}$ be the time after which the bounds in Lemma 7 and Lemma 8 hold. Then for $t>T_{0}$,
$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}]$
$\leq 2\mathbb{E}[\|\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))\|^{2}]+2\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]$
$\leq 2\|\boldsymbol{A}\|^{2}\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+2\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]$
$\leq C\left(\eta_{t}+1/(t\epsilon_{t})\right),$
for $C=\max\{2\|\boldsymbol{A}\|^{2}C^{\prime},2C^{\prime\prime}\}$.